OpenStudy (anonymous):

Show me how to do this step by step. Let me put the equation on the 2nd post.

OpenStudy (king):

equation?

OpenStudy (anonymous):

\[\sqrt{5x+4}-1=\sqrt{3x+1}\]

OpenStudy (mr.math):

Square both sides of the equation \[5x+4-2\sqrt{5x+4}+1=3x+1 \implies 2x+3=2\sqrt{5x+4}\] Square both sides of the equation again \[4x^2+12x+9=4(5x+4)\] This is a quadratic equation now, I think you can do it.

OpenStudy (king):

hmmm....i got the quad. equation as 4x^2-20x+8=0... and so x shud be (5+sqrt17)/2, (5-sqrt17)/2

OpenStudy (anonymous):

hm okay

OpenStudy (anonymous):

anymore hints? i'm not getting this at all.

OpenStudy (mr.math):

I made a slight mistake above. We had \[5x+4-2\sqrt{5x+4}+1=3x+1\implies 2x+4=2\sqrt{5x+4}.\] Squaring both sides gives: \[4x^2+16x+16=4(5x+4) \implies 4x^2+16x+16=20x+16 \] \[\implies 4x^2-4x=0 \implies 4x(x-1)=0 \implies x=0 \text{ or }x=1.\] Now you need to check each solution by plugging into the original equation. At x=0, we get 1=1, which is true. At x=1, we have 2=2. so x=1 is also a solution.